Suppose atomispheric CO2 is growing at 2ppm/yr, fossil fuel and cement emissions are 9GtC/yr, and the airborne fraction is 38%. If the only other carbon emissions are associated with land use changes, estimate their net carbon emission rate.
Answer #1

Given,

Fossil fuel and cement emission = 9 GtC/yr

CO2 level = 2 ppm/yr = 2 * 2.12 = 4.24 GtC/yr

Land use level = 2.2 GtC/yr

Airborne fraction = 38%

Total Concentration = Fossil fuel and cement emission + CO2 level + Land use level = 9 + 4.24 + 2.2 = 15.44 GtC/yr

Net Carbon Emission Rate = (15.44 * 0.38)/2.12 = 2.7675 ppm/yr CO2
Answer #2

Suppose global Fossil fuel combustion emists 7.7 GTC/yr and cement production add 0.5 GTC. Assume air born fraction as 0.38 and assuming no change in emissions due to land use, what rate of change co2 concentration would you expect